OpenStudy (anonymous):

obtain Yn for y=e^ax . cos(bx + c) and find y- e^x . cos 5x . cos 3x

OpenStudy (anonymous):

hint cos 5x . cos 3x =1/2 * (cos(5+3)x + cos(5-3)x)=1/2 * (cos8x +cos2x)